OpenStudy (anonymous):

In order to solve for the variable in the equation 2(x+3)+5x=3(2x-1) , Jaleesa begins by applying the distributive property, then combines like terms. Which equation is the result of these steps?

OpenStudy (anonymous):

2(x+3)+5x=3(2x-1) distribute 2x+6+5x=6x-3 combine 7x+6=6x-3